# Spatial-aware Transformer-GRU Framework for Enhanced Glaucoma Diagnosis from 3D OCT Imaging

### v1

@@ 4.4 Ablation Study
Table 5: Summary of ablation study results averaged from a 5-fold cross-validation with 95% confidence intervals (±).
Method ACC AUC SEN SPE PRC F1-score MCC
ResNet34 + GRU
87.39 (± 1.73) 92.99 (± 1.75) 92.03 72.86 91.38 91.71 65.40
ViT-large + SVM (Slice 1)
77.47 (± 3.39) 83.79 (± 1.25) 91.29 51.68 77.92 84.08 48.15
ViT-large + SVM (Slice 31)
80.09 (± 1.43) 88.08 (± 1.17) 92.18 55.71 80.76 86.09 53.02
ViT-large + SVM (Slice 32)
82.44 (± 2.69) 89.34 (± 1.46) 92.89 59.71 83.35 87.87 57.49
ViT-large + SVM (Slice 37)
82.52 (± 3.58) 89.34 (± 2.23) 92.79 59.94 83.59 87.95 57.49
ViT-large + SVM (Slice 52)
78.11 (± 3.71) 86.25 (± 1.98) 92.42 52.51 77.69 84.41 50.67
ViT-large + SVM (Majority voting)
85.05 (± 2.23) 92.62 (± 1.33) 94.16 64.31 85.71 89.74 63.33
Proposed (ViT-large + GRU)
90.27 (± 1.44) 95.24 (± 1.69) 94.25 78.18 92.92 93.58 73.54
@@
In the subsequent experiment, we assessed the impact of replacing the GRU model in the proposed framework with a voting ensemble model. To address the challenge of incorporating all available slices in 3D OCT images, we initially reduced the number of slices. Therefore, from the original set of 64 OCT slices, we chose 5 slices based on their entropy values. Entropy, in this context, measures the information content for identifying the most relevant slices for glaucoma detection. We computed the entropy of all slices in the OCT images. Considering the significance of the central slice in glaucoma diagnosis [33], we selected four slices with entropy values close to the $32^{\mathrm{nd}}$ slice and the central slice itself. This method ensures a balance between data-driven selection and clinical relevance.
@@ 5 Conclusion
The experimental results validate the framework’s superior performance, achieving an AUC of 95.24%, F1-score of 93.58%, and MCC of 73.54%. These metrics outperform state-of-the-art methods, including the 3D-CNN [20] and RETFound model [35] extended for slice-based classification. These results underscore the potential of the proposed approach to address challenges posed by class imbalance and subtle glaucomatous patterns distributed across the ocular anatomy, enhancing clinical decision support systems and improving patient outcomes in glaucoma management.
In our opinion, there are several opportunities for future research. Exploring the integration of complementary clinical data modalities, such as visual field tests or patient demographics, could further enhance the diagnostic capabilities of the proposed framework. Moreover, investigating alternative sequence processing models and incorporating attention mechanisms may yield further performance improvements in glaucoma detection and screening tasks. Additionally, extending the application scope of our framework coupled with interpretability to other ocular pathologies could significantly contribute to the advancement of early diagnosis and treatment strategies within the broader field of ophthalmology. Furthermore, developing robust techniques to mitigate the impact of data imbalance could substantially enhance the diagnostic accuracy and generalizability of the framework across diverse datasets and imaging modalities. Such strategies would enable a comprehensive assessment of the framework’s robustness and facilitate its broader deployment in clinical ophthalmological practice.

### v2

@@ 4.4 Ablation Study
Table 5: Summary of ablation study results averaged from a 5-fold cross-validation with 95% confidence intervals (±). Values are reported as percentages, with the confusion matrix relative to the total samples per class.
Method ACC AUC SEN SPE PRC F1-score MCC Confusion Matrix
ResNet34 + GRU 87.21 (± 2.22) 93.55 (± 2.53) 90.85 72.86 92.56 91.70 63.92 92.56 7.44
30.04 69.96
ViT-large + LSTM 87.51 (± 2.88) 93.63 (± 2.64) 91.65 73.85 91.97 91.81 65.23 91.97 8.03
27.00 73.00
ViT-large + SVM (Slice 1) 77.32 (± 2.95) 83.75 (± 2.35) 91.38 51.41 77.57 83.91 48.06 77.57 22.43
23.58 76.42
ViT-large + SVM (Slice 31) 78.46 (± 4.46) 85.55 (± 4.03) 91.19 53.23 79.46 84.92 49.31 79.47 20.53
24.71 75.29
ViT-large + SVM (Slice 32) 79.42 (± 3.86) 87.30 (± 2.83) 91.54 54.79 80.52 85.68 51.20 80.53 19.47
23.95 76.05
ViT-large + SVM (Slice 37) 79.10 (± 1.41) 86.58 (± 3.40) 91.61 54.18 79.93 85.37 50.80 79.93 20.07
23.58 76.42
ViT-large + SVM (Slice 52) 76.77 (± 2.77) 85.37 (± 2.88) 90.62 50.65 77.57 83.59 46.20 77.57 22.43
25.86 74.14
ViT-large + SVM (Majority voting) 83.14 (± 1.87) 91.55 (± 2.10) 93.19 60.98 84.06 88.39 59.02 84.05 15.95
19.77 80.23
Proposed (ViT-large + GRU) 89.19 (± 1.89) 94.20 (± 2.00) 91.83 79.67 94.21 93.01 69.33 94.21 5.79
26.98 73.02
@@
To further investigate the effect of different RNN layers on sequential dependency modeling, we replaced the GRU layers with LSTM layers. The LSTM employs a more complex gating mechanism to capture long-term dependencies more effectively. However, as reported in Table 5, the GRU layers marginally outperformed the LSTM layers. This result suggests that the simpler gating mechanism of the GRU reduces model complexity, making it more suitable for our framework. In the subsequent experiment, we assessed the impact of replacing the sequential processing in the proposed framework with a voting ensemble model. To address the challenge of incorporating all available slices in 3D OCT images, we initially reduced the number of slices. Therefore, from the original set of 64 OCT slices, we chose 5 slices based on their entropy values. Entropy, in this context, measures the information content for identifying the most relevant slices for glaucoma detection. We computed the entropy of all slices in the OCT images. Considering the significance of the central slice in glaucoma diagnosis [34], we selected four slices with entropy values close to the $32^{\mathrm{nd}}$ slice and the central slice itself. This method ensures a balance between data-driven selection and clinical relevance.
@@ 5 Conclusion
The experimental results validate the framework’s superior performance, achieving an AUC of 94.20%, F1-score of 93.01%, and MCC of 69.33%. These metrics outperform state-of-the-art methods, including the 3D-CNN [21] and RETFound model [36] extended for slice-based classification. These results underscore the potential of the proposed approach to address challenges posed by class imbalance and subtle glaucomatous patterns distributed across the ocular anatomy, enhancing clinical decision support systems and improving patient outcomes in glaucoma management.
In our opinion, there are several opportunities for future research. Exploring the integration of complementary clinical data modalities, such as visual field tests or patient demographics, could further enhance the diagnostic capabilities of the proposed framework. Moreover, investigating alternative sequence processing models and incorporating attention mechanisms may yield further performance improvements in glaucoma detection and screening tasks. Additionally, extending the application scope of our framework coupled with interpretability to other ocular pathologies could significantly contribute to the advancement of early diagnosis and treatment strategies within the broader field of ophthalmology. Furthermore, developing robust techniques to mitigate the impact of data imbalance could substantially enhance the diagnostic accuracy and generalizability of the framework across diverse datasets and imaging modalities. However, it is important to note that deploying this framework in clinical settings would require rigorous clinical trials, expert validation, and integration into real-world workflows that involve patients from diverse ethnicities, genders, and other demographic groups. Such strategies would enable a comprehensive assessment of the framework’s robustness and facilitate its broader deployment in clinical ophthalmological practice.